A particle is moving in a one dimensional width 25 Armstrong assuming the particle is in its least energy state calculate the probability of finding particle in an interval of 5 Armstrong at the distances of A/2,A/3 and at A where a is the width of the potential well


1
Expert's answer
2022-04-04T09:04:00-0400

Answer

Width l=10"A\u00b0"

Interval

"x=5A\u00b0"

Since the electron is in its least energy state,

n=1

So probabilty is

"\psi_n=\sqrt{\frac{2}{l}}sin(\frac{n\pi x}{l})"

Then probabilty

"\psi_n=\sqrt{\frac{2}{10}}sin(\frac{\pi 5}{10}) \\=0.45"